The odds that an inspector correctly identifies a damaged sample in a textile factory are estimated to be 95%. The odds that a sample is incorrectly identified as damaged is 2%. Moreover the percentage of damaged samples produced by this factory is 1%.

a) What is the probability that a sample selected for inspection is identified as damaged?

b) If a sample chosen randomly is identified as damaged, what is the probability that the sample is actually NOT damaged?
Answer #1

a)  probability that a sample selected for inspection is identified as damage = P(correct identification and damage) + P(incorrect identification and undamage) = 0.01*0.95 + 0.02*(1-0.01)= 0.0293 or 2.93%

b)

P(actually not damage | identified as damage) =  0.02*(1-0.01)/0.0293 = 0.6758 or 67.58%